A high school sold gift wrap. The gift wrap in solid colors sold for $4.00 per roll and the print gift wrap sold for $6.00 per roll. The total amount of rolls sold was 480, and the total amount of money collected was $2,340.00. How many rolls of each kind of gift wrap was sold?
Answer by nerdybill(5413)

A high school sold gift wrap. The gift wrap in solid colors sold for $4.00 per roll and the print gift wrap sold for $6.00 per roll. The total amount of rolls sold was 480, and the total amount of money collected was $2,340.00. How many rolls of each kind of gift wrap was sold?
.
Let x = number of $4 rolls sold
then
480-x = number of $6 rolls sold
.
4x + 6(480-x) = 2340
4x + 2880 - 6x = 2340
2880 - 2x = 2340
2880 = 2x + 2340
540 = 2x
270 = x ($4 rolls sold)
.
480-x = 480-270 = 210 ($6 rolls sold)